Franklin Construction entered into a fixed-price contract to build a freeway-connecting ramp for $30 million. Construction costs incurred in the first year were $16 million and estimated remaining costs to complete at the end of the year were $17 million.

How much gross profit or loss will Franklin recognize in the first year if it recognizes revenue over time according to percentage of completion method?

How much gross profit or loss will Franklin recognize in the first year applying the completed contract method?
Answer #1

Answer A

Total Contract price =$30

Total cost incurred =$16

Estimated remaining cost =$17

Total estimated cost of contract =$33

Percentage of completion = cost incurred/total cost estimated×100

= 16/33×100 = 48.49%

Profit/Loss for 1st year

Loss as per percentage of completion method = $30×48.49%= $14.545

Answer b

The completed contract method is used to recognize all of the revenue and profit associated with a project only after the project has been completed. This method is used when there is uncertainty about the collection of funds due from a customer under the terms of a contract. This method yields the same results as the percentage of completion method, but only after a project has been completed. Prior to completion, this method does not yield any useful information for the reader of a company’s financial statements. However, the delay in income recognition allows a business to defer the recognition of related income taxes.

Also, since revenue and expense recognition only occurs at the end of a project, the timing of revenue recognition can be both delayed and highly irregular. Given these issues, the method should only be used under the following circumstances:

  • When it is not possible to derive dependable estimates about the percentage of completion of a project; or

  • When there are inherent hazards that may interfere with completion of a project; or

  • When contracts are of such a short-term nature that the results reported under the completed contract method and the percentage of completion method would not vary materially.

If a contract is being accounted for under this method, record billings issued and costs incurred on the balance sheet during all periods prior to the completion of the contract, and then shift the entire amount of these billings and costs to the income statement upon completion of the underlying contract. A contract is assumed to be complete when the remaining costs and risks are insignificant.

If there is an expectation of a loss on a contract, record it at once even under the completed contract method; do not wait until the end of the contract period to do so.

Hence expected loss = total estimated cost - contract price

=$33-$30= $3 (expected loss)